OpenStudy (anonymous):

write the percent as a fraction or a mixed number in simplest form 4 2/5%. Can anyone help me solve this?

OpenStudy (anonymous):

i have a picture to better understand it

OpenStudy (ranga):

4 & 2/5 = (5*4 + 2)/5 = 22/5%. Convert percentage to a fraction by dividing by 100: 22/5% = 22/500 = 11/250. So the simplest fraction is: 11/250.

OpenStudy (anonymous):

omg thank you thank you soo much

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

and is that how i figure out all problems like this

OpenStudy (ranga):

Yes. First you convert the mixed number into a fraction. Examples:\[4\frac{ 2 }{ 5 } = \frac{ 5*4 + 2 }{ 5 } = \frac{ 22 }{ 5 }\]\[12\frac{ 5 }{ 9 } = \frac{ 9*12 + 5 }{ 9 } = \frac{ 113 }{ 9 }\]Here you have to do one extra step. The original mixed number was a percentage. To convert percentage to a fraction you will have to divide by 100. Therefore,\[\frac{ 22 }{ 5 }\% = \frac{\frac{ 22 }{ 5 }}{100} = \frac{22}{500} = \frac{11}{250}\]\[\frac{ 113 }{ 9 }\% = \frac{\frac{ 113 }{ 9 }}{100} = \frac{113}{900}\]